Determine approximately the ratio of the probability of spontaneous emission


to the probability of stimulated emission at room temperature in (a) X-ray


region of the EM spectrum (b) the visible region (c) the microwave region.

1
Expert's answer
2021-12-17T18:16:08-0500

"u=\frac{8\pi f^2}{c^3}\cdot \frac{hf}{e^{\frac{hf}{kT}-1}},"

"u_1=1.7\cdot 10^{-4},~~\\"

"u_2=5.1\cdot 10^{-5},"

"u_3=2.0\cdot 10^{-7}."